OpenStudy (anonymous):

Identify the solutions of 3x2 - x - 4 = 0. Answer

OpenStudy (anonymous):

that is the quadratic equation... do you know its formula ? :)

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

\[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\LARGE ax^2+bx+c=0\] in your case you have: a=3 b=-1 c=-4 substitute in this formula: \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and there you go... :) can you do it ?

OpenStudy (anonymous):

yep got it thanks

OpenStudy (anonymous):

My pleasure :)